OpenStudy (anonymous):

evaluate each equation and solve each eqution. 1. x+8=13 2.x/4= -32 3. 16+2n=10 4. 3(h+w);h = -5 and w=4 write an algebric expretion for each of the given phrase. 5. 3 inches less than 2 times the length x of the given table in inches. 6. the quotent of 25 and a number p.

OpenStudy (merchandize):

1) x+8=13 x=13-8 x=5

OpenStudy (anonymous):

1.) x + 8 = 13 First subtract 8 from both sides. -8 -8 -------------- x = 5 2.) x/4= -32 You have to multiply -32 and 4 to get x x = -128 3.) 16+2n=10 You first have to subtract 16 from both sides -16 -16 ----------------- 2n = -6 Then divide both sides by 2. /2 /2 -------- n = -3 4.) 3(h+w);h = -5 and w=4 Youave first have to substitute h for -5 and w for fou 3(-5 + 4) Then distribute -15 + 12 Then add -3

OpenStudy (merchandize):

4) 3(h+w) put values of h and w 3(-5+4) =3(-1) -3

OpenStudy (anonymous):

5. 2x - 3 6. 25/p

OpenStudy (anonymous):

thank u!!!!!